Circular No. 5647

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

GX 339–4 B. A. Harmon, G. J. Fishman, W. S. Paciesas, and M. Finger, Mar-

shall Space Flight Center, NASA, report for the Burst and Transient Source Experiment (BATSE)/Compton Observatory Team: “The black-hole candidate GX 339–4 has entered its hard (low) emission state, being detected by earth occultation beginning about mid-September. Its intensity has increased very gradually and is currently at ∼ 0.24 Crab (Oct. 17-25) in the 20- to 300-keV band. The spectrum is very similar to the 1991 hard-state outburst (IAUC 5327), extending to at least 230 keV, with a power law index of –2.1 ± 0.06. A better fit is obtained, however, with a Comptonization model, as the spectrum becomes steeper above 100 keV. The Sunyaev-Titarchuk model fit parameters for Oct. 17-25 are kT = 39 ± 5 keV and γ = 2.5 ± 0.5. Identification of GX 339–4 as the flaring source was obtained by fitting for the time of occultation edges, yielding a 1σ error in location of ± 0^{o}.35 for α and ± 0^{o}.2 for δ. Monitoring of the source intensity and spectrum is continuing.”

JUPITER G. Orton, P. Yanamandra-Fisher, and D. Griep, NASA Infrared Tele-

scope Facility, report: “Eighteen-μm images of Jupiter on Oct. 26.68-26.78 UT show that the South Equatorial Belt (SEB) has resumed a 150-mbar temperature equal to that of the NEB, recovering from prior cooler temperatures which lasted ∼ 1 yr. Infrared indicators at 8.57 and 4.9 μm show the SEB remaining unusually cloudy, as it has been since early 1992. If it behaves as in 1988-1990, the SEB warming signals the recovery of its cloud properties to a more typical state within 6-9 months.”

NOVA SAGITTARII 1992 No. 2 Further photometry (cf. IAUC 5633) by A. C. Gilmore, Mount John

University Observatory: Oct. 17.42 UT, V = 12.62 ± 0.08, B–V = +0.21 ± 0.08, U–B = –0.26 ± 0.08, V–R = +1.12 ± 0.09.

NOVA SAGITTARII 1992 No. 3 Further photometry (cf. IAUC 5642) by Gilmore: Oct. 24.38 UT, V =

8.92 ± 0.04, B–V = +0.40 ± 0.03, U–B = –0.40 ± 0.05, V–R = +0.56 ± 0.03, V–I = +0.87 ± 0.03.

1992 November 3 Daniel W. E. Green